OpenStudy (anonymous):

In how many ways can you roll either a sum of 4 or a sum of 11 with a pair of dices?

OpenStudy (anonymous):

It says the answer is 5,

random231 (random231):

@M0j0jojo are u trying to do a probability sum?

OpenStudy (anonymous):

no this permutations

random231 (random231):

yeah 5

OpenStudy (anonymous):

can you explain please

OpenStudy (anonymous):

Assuming the dices are not thrown at the same time, let A: Result of the first dice B: Result of the second dice A={1,2,3,4,5,6} B={1,2,3,4,5,6} A+B=4 (Condition to roll a sum of four) 1+3=4 2+2=4 2+2=4 3+1=4 (Four ways to roll a sum of four if you care about the order of the dices, but only two ways if you throw the dices at the same time and only care about the resulting sum) A+B=11 (Condition to roll a sum of eleven) 5+6=11 6+5=11 (Two ways to roll a sum of four if you care about the order of the dices, but only one way if you throw the dices at the same time and only care about the resulting sum) I think there must be a way to solve this kind of problems more generically using combinatorics, but I don't quite remember it, I'll see if I can find something about that.

random231 (random231):

well 2 2 1 3 3 1 6 5 5 6 u hav to take all the combinations

OpenStudy (destinymasha):

actually, it is 5 -- my bad

random231 (random231):

yeah u hav to care the order of the dices

OpenStudy (anonymous):

oh okay i get it

OpenStudy (anonymous):

thanks guys